# Simulation-based inference of galaxy properties from JWST pixels

### v1

@@ Acknowledgements.
Co-funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them. PIN thanks the LSST-DA Data Science Fellowship Program, which is funded by LSST-DA, the Brinson Foundation, the WoodNext Foundation, and the Research Corporation for Science Advancement Foundation; her participation in the program has benefited this work. MHC and PIN acknowledge financial support from the State Research Agency of the Spanish Ministry of Science and Innovation (AEI-MCINN) under the grants “Galaxy Evolution with Artificial Intelligence” with reference PGC2018-100852-A-I00 and “BASALT” with reference PID2021-126838NB-I00. JHK acknowledges support from AEI-MCINN under the grant “The structure and evolution of galaxies and their outer regions” and the European Regional Development Fund (ERDF) with reference PID2022-136505NB-I00/10.13039/501100011033. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa.
{dataavailability}
@@ Appendix A Posterior distributions using different SFH priors for a simulated galaxy
We include in Fig. 8 two corner plots showing the posterior distributions obtained for a single simulated galaxy (using a $\tau$-delayed simulation prior) when inference is performed with the model trained on the $\tau$-delayed simulation (left) and the model trained on the Dirichlet simulation (right). The posterior distributions are almost identical for the two models, although the distributions are slightly wider for the model trained on the Dirichlet prior. We can also see the correlations and anticorrelations that exist between the different properties. On the left we find positive correlations between $\log_{10}(M_{*}/\rm{M}_{\odot})$ and $\log_{10}$(SFR/($\rm{M}_{\odot}$ yr-1)), $\log_{10}$(SFR/($\rm{M}_{\odot}$ yr-1)) and $A_{V}$, and $t_{i}$ (cosmic time) and [$M$/H], as well as negative correlations between $\log_{10}(M_{*}/\rm{M}_{\odot})$ and [$M$/H], and [$M$/H] and $A_{V}$. On the right, these are less obvious because the distributions are wider, but we observe positive correlations between different $t_{x}$ (cosmic time normalised to the age of the Universe at the redshift of the source), and negative correlations between [$M$/H] and $A_{V}$.
@@ Appendix B Simulation-based calibration
We perform a simulation-based calibration (SBC; Talts et al. 2018) test to ensure that the NFs are well trained and thus the posteriors are properly calibrated. In Fig. 9 we show the distribution of the rank statistics of the true parameter values within the marginalised posteriors for the $\tau$-delayed model, which must be uniform if the posterior samples are consistent with the prior used in the simulations. We show in Fig. 10 the analogue for the Dirichlet prior. Figure  11 shows the empirical cumulative distribution function for both the $\tau$-delayed model and the Dirichlet model, which confirms that they produce posteriors that are fully consistent with the simulations, without underestimating or overestimating the uncertainties. However, we point out that the performance of any model when obtaining posteriors distributions for real galaxies is unknown, and the ability to get accurate distributions, as in any inference pipeline, ultimately depends only on how well the forward model (or likelihood in traditional approaches) matches the observations.
@@ Appendix C Posterior distributions as a function of the S/N
Both the surviving and formed stellar masses can be well-constrained up to 15% of the prior distribution width, without important changes with the S/N, as low S/N are also informative of the mass content of each of pixels up to S/N $\sim 1$. There are no strong differences with redshift either, with only a slight increment across all S/N. The dust attenuation index $A_{V}$ can also be recovered with narrow posterior distributions ($\sigma\sim 0.2\sigma_{\rm{prior}}$), but experiment a sudden broadening at S/N $<5$, meaning that we start to be prior-dominated. Interestingly, $A_{V}$ can be better constrained at higher redshifts, probably because most galaxies are dust free, but also because we have access to the restframe UV slope and Balmer break for these galaxies.
@@
The behaviour of the rest of the properties depends on the prior used for the SFHs. The metallicity cannot be retrieved with precision, as already known from photometry only (e.g., Nersesian et al. 2025), we only reach $\sigma\sim 0.6\sigma_{\rm{prior}}$ for the $\tau$-delayed prior, and $\sigma\sim 0.7\sigma_{\rm{prior}}$ for the Dirichlet prior, experimenting also a significant broadening for the posteriors at S/N $<5$. While for the Dirichlet prior we do not find important changes across redshifts, for the $\tau$-delayed prior we find slightly narrower posteriors at high redshift, maybe highlighting a strong degeneracy with the age that is only significant at low redshift due to the larger possible ages.
@@
The SFR behaviour is very different from one prior to another, with $\sigma\sim 0.15\sigma_{\rm{prior}}$ and $\sigma\sim 0.4\sigma_{\rm{prior}}$ at low redshifts respectively, as expected by restricting the shapes of the SFHs by using a functional form. We obtain overall narrower distributions at high-redshift, especially for the Dirichlet prior. This can again be due to lower dust attenuation but also due to selection effects at higher redshifts in the ranges of masses analysed. The parameter $t_{i}$ can be poorly constrained, with values ranging from $\sigma\sim 0.6-0.8\sigma_{\rm{prior}}$, slightly better for $z>5$. For $\tau$, $t_{25\%}$, $t_{50\%}$ and $t_{75\%}$, we should not compare the values with the other properties because the metric is different. However, we find strong broadening of the posterior distributions for S/N $<5$. The width of the $\tau$ posterior distributions is higher with redshift, but slightly lower for $t_{25\%}$ and $t_{75\%}$. As expected, in most of the redshift bins $t_{25\%}$ is wider than $t_{75\%}$, but this result is inverted for the last bin of redshift. We expect this to be a consequence of the prior for $t_{25\%}$ that highly favours short cosmic times. For the parametric prior, we also combine the $\tau$ and $t_{i}$ samples of the posteriors to derive samples for the mass-weighted age, which is known to be a more robust and less degenerated quantity, and find that we can indeed constrain better this property with $\sigma\sim 0.4\sigma_{\rm{prior}}$ for all redshift bins.
@@ Appendix E Recovered SEDs: posterior predictive checks
We perform a PPC (Talts et al. 2018), repeating the simulation for every sample of the posteriors for the same two pixels of each of the six galaxies using the parametric model. We compare the results of the PPC with the actual observed photometry in Fig. 14. We find a general agreement across all galaxies. However, certain sets of SEDs deviate from the true photometry in some filters, typically the bluest ones, for instance F775W and F814W for the pixel with highest S/N of 206146. This discrepancy arises from model misspecification when none of the simulation templates align with the observed photometry for any noise realization, making the posterior distributions unreliable. This misalignment could stem from problems in the photometry of that pixel, such as contamination from other objects or improper background subtraction. However, it is more likely due to assumptions in the SPS models, including the prior ranges. On the other hand, we observe for some of the pixels there multiple possible SED templates and the model correctly interprets this, so that the posterior distributions are wider, especially at low S/N and at high redshifts where several filters are only fitted with upper limits. The best example for this is the pixel outside $1R_{\rm{eff}}$ of 211273.
We include in the same figure the best fit of Prospector in pink, as well as the Synthesizer fit in green. While the fit of Prospector uses the same SPS modelling assumptions, and the deviations we see with respect of our model are mainly random, the differences with Synthesizer are systematic, as we can clearly see different spectral templates. This is more evident for high-redshift galaxies and show the importance of the choice of libraries and parameters for stellar population properties estimation and of marginalizing over all the different templates to properly capture the true uncertainties.
@@ Appendix F Profiles of stellar population properties
We compute profiles of stellar mass density, mass-weighted age and dust attenuation, from posterior distributions obtained with the parametric model. In Fig. 15, we show in blue the profiles of the properties computed from elliptical apertures with the semimajor axis, the ellipticity and the position angle obtained from the JADES photometry catalogue (Eisenstein et al. 2023a), and the centres fixed as the pixel with higher stellar masses (medians of the posteriors). The effective radius we use to normalise the distances to the centre was also obtained from the JADES catalogue. The profiles correspond to the medians of the posterior distributions and the errors to the $1\sigma$ level, normalized to the values in the centre of the galaxy. It is important to note again that the profiles below the PSF FWHM of F [^4] W are mainly flat, and after this radial scale vary significantly.
We recover what we see in the previous maps, the stellar mass density profiles are smooth and do not suffer strong variations, while the other properties, especifically for the first three galaxies where we reach larger radius, show important variations. While the uncertainties ($1\sigma$) are considerable, we find similar profiles for both three low redshift galaxies (first row), with medium-age centres, old regions between $0.2-0.5$ $R_{\rm{eff}}$ and younger outskirts. For the high-redshift galaxies, we obtain flatter profiles, specially for 118081 where we are mostly inside the FHWM of F [^4] W due to the compactness of the galaxy. For the first four galaxies the centres are more dust attenuated than the outskirts, but both 118081 and 206146 have very little dust contribution in the centre, compatible with the high redshift of these sources.
@@ Appendix G Offset in stellar mass estimations with different priors for the SFHs
For the $\tau$-delayed model applied to the Dirichlet simulation we observe a significant average underestimation of $0.15$ dex. This is expected because the fixed form of the $\tau$-delayed prior is not able to capture early episodes of star formation (e.g., Carnall et al. 2019; Lower et al. 2020). We also compute the fraction of the true total stellar mass formed before the median $t_{i}$ obtained with the $\tau$-delayed model, according to the simulated Dirichlet SFH. We find a correlation between the residuals and this fraction, showing that the largest underestimations of total stellar mass for the $\tau$-delayed model actually occur when there are episodes of star formation that contribute significantly to the mass, but are obscured by more moderate and recent bursts of star formation.
@@
In the second row, the simulation is performed with a $\tau$-delayed prior. While in the left case, where $\tau$ is large and thus the SFH is extended, the Dirichlet model performs well and also obtains a recent SFR that is close to the true value, in the right case the Dirichlet model predicts previous episodes of star formation that are not there, leading to $69\%$ of the mass predicted to form before the star formation actually starts. Nevertheless, in this case we observe that some of the SFHs drawn from the posterior samples are fully compatible with the simulation. This flexibility of the Dirichlet prior, perhaps excessive, allows us to marginalise through a complete set of behaviours, and in this case the large uncertainties in constraining the SFHs are present in the posterior distributions of all the parameters. This leads to broader posterior distributions, as already shown in Fig. 12, which are more representative of the ignorance of the past SFHs of true galaxies. This effect is also found in the pixel-by-pixel analysis of true JWST galaxies. While for most galaxies the flexibility of the pixel-by-pixel analysis probably does not require complex SFHs at the pixel level, one pixel may represent several SSPs with non-trivial SFHs, especially at high redshift. Ultimately, we should look for physically motivated priors that match the burstiness and star formation scales of the real Universe. This approach, already started by the works of i.e., Caplar & Tacchella (2019); Tacchella et al. (2020); Iyer et al. (2024); Wan et al. (2024) and Wang et al. (2025), is crucial to properly assess the estimation of stellar population parameters with well-calibrated uncertainties from photometry alone.

### v3

@@ Acknowledgements.
Co-funded by the European Union (MSCA Doctoral Network EDUCADO, GA 101119830 and Widening Participation, ExGal-Twin, GA 101158446). PIN thanks the LSST-DA Data Science Fellowship Program, which is funded by LSST-DA, the Brinson Foundation, the WoodNext Foundation, and the Research Corporation for Science Advancement Foundation; her participation in the program has benefited this work. MHC and PIN acknowledge financial support from the State Research Agency of the Spanish Ministry of Science and Innovation (AEI-MCINN) under the grants “Galaxy Evolution with Artificial Intelligence” with reference PGC2018-100852-A-I00 and “BASALT” with reference PID2021-126838NB-I00. JHK acknowledges support from AEI-MCINN under the grant “The structure and evolution of galaxies and their outer regions” and the European Regional Development Fund (ERDF) with reference PID2022-136505NB-I00/10.13039/501100011033. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa.
@@ Appendix A Posterior distributions using different SFH priors for a simulated galaxy
We include in Fig. 8 two corner plots showing the posterior distributions obtained for a single simulated galaxy (using a $\tau$-delayed simulation prior) when inference is performed with the model trained on the $\tau$-delayed simulation (left) and the model trained on the Dirichlet simulation (right). The posterior distributions are almost identical for the two models, although the distributions are slightly wider for the model trained on the Dirichlet prior. We can also see the correlations and anticorrelations that exist between the different properties. On the left we find positive correlations between $\log_{10}(M_{*}/\rm{M}_{\odot})$ and $\log_{10}$(SFR/($\rm{M}_{\odot}$ yr-1)), $\log_{10}$(SFR/($\rm{M}_{\odot}$ yr-1)) and $A_{V}$, and $t_{i}$ (cosmic time) and [MM/H], as well as negative correlations between $\log_{10}(M_{*}/\rm{M}_{\odot})$ and [MM/H], and [MM/H] and $A_{V}$. On the right, these are less obvious because the distributions are wider, but we observe positive correlations between different $t_{x}$ (cosmic time normalised to the age of the Universe at the redshift of the source), and negative correlations between [MM/H] and $A_{V}$.
@@ Appendix B Simulation-based calibration
We perform a simulation-based calibration (SBC; Talts et al. 2018) test to ensure that the NFs are well trained and thus the posteriors are properly calibrated. In Fig. 9 we show the distribution of the rank statistics of the true parameter values within the marginalised posteriors for the $\tau$-delayed model, which must be uniform if the posterior samples are consistent with the prior used in the simulations. We show in Fig. 10 the analogue for the Dirichlet prior. Figure 11 shows the empirical cumulative distribution function for both the $\tau$-delayed model and the Dirichlet model, which confirms that they produce posteriors that are fully consistent with the simulations, without underestimating or overestimating the uncertainties. However, we point out that the performance of any model when obtaining posteriors distributions for real galaxies is unknown, and the ability to get accurate distributions, as in any inference pipeline, ultimately depends only on how well the forward model (or likelihood in traditional approaches) matches the observations.
@@ Appendix C Posterior distributions as a function of the S/N
Both the surviving and formed stellar masses can be well-constrained up to 15% of the prior distribution width, without important changes with the S/N, as low S/N are also informative of the mass content of each of pixels up to S/N $\sim 1$. There are no strong differences with redshift either, with only a slight increment across all S/N. The dust attenuation index $A_{V}$ can also be recovered with narrow posterior distributions ($\sigma\sim 0.2\sigma_{\rm{prior}}$), but experiment a sudden broadening at S/N <5<5, meaning that we start to be prior-dominated. Interestingly, $A_{V}$ can be better constrained at higher redshifts, probably because most galaxies are dust free, but also because we have access to the restframe UV slope and Balmer break for these galaxies.
@@
The behaviour of the rest of the properties depends on the prior used for the SFHs. The metallicity cannot be retrieved with precision, as already known from photometry only (e.g. Nersesian et al. 2025), we only reach $\sigma\sim 0.6\sigma_{\rm{prior}}$ for the $\tau$-delayed prior, and $\sigma\sim 0.7\sigma_{\rm{prior}}$ for the Dirichlet prior, experimenting also a significant broadening for the posteriors at S/N <5<5. While for the Dirichlet prior we do not find important changes across redshifts, for the $\tau$-delayed prior we find slightly narrower posteriors at high redshift, maybe highlighting a strong degeneracy with the age that is only significant at low redshift due to the larger possible ages.
@@
The SFR behaviour is very different from one prior to another, with $\sigma\sim 0.15\sigma_{\rm{prior}}$ and $\sigma\sim 0.4\sigma_{\rm{prior}}$ at low redshifts respectively, as expected by restricting the shapes of the SFHs by using a functional form. We obtain overall narrower distributions at high-redshift, especially for the Dirichlet prior. This can again be due to lower dust attenuation but also due to selection effects at higher redshifts in the ranges of masses analysed. The parameter $t_{i}$ can be poorly constrained, with values ranging from $\sigma\sim 0.6-0.8\sigma_{\rm{prior}}$, slightly better for $z>5$. For $\tau$, $t_{25\%}$, $t_{50\%}$ and $t_{75\%}$, we should not compare the values with the other properties because the metric is different. However, we find strong broadening of the posterior distributions for S/N <5<5. The width of the $\tau$ posterior distributions is higher with redshift, but slightly lower for $t_{25\%}$ and $t_{75\%}$. As expected, in most of the redshift bins $t_{25\%}$ is wider than $t_{75\%}$, but this result is inverted for the last bin of redshift. We expect this to be a consequence of the prior for $t_{25\%}$ that highly favours short cosmic times. For the parametric prior, we also combine the $\tau$ and $t_{i}$ samples of the posteriors to derive samples for the mass-weighted age, which is known to be a more robust and less degenerated quantity, and find that we can indeed constrain better this property with $\sigma\sim 0.4\sigma_{\rm{prior}}$ for all redshift bins.
@@ Appendix E Recovered SEDs: posterior predictive checks
We perform a PPC (Talts et al. 2018), repeating the simulation for every sample of the posteriors for the same two pixels of each of the six galaxies using the parametric model. We compare the results of the PPC with the actual observed photometry in Fig. 14. We find a general agreement across all galaxies. However, certain sets of SEDs deviate from the true photometry in some filters, typically the bluest ones, for instance F775W and F814W for the pixel with highest S/N of 206146. This discrepancy arises from model misspecification when none of the simulation templates align with the observed photometry for any noise realisation, making the posterior distributions unreliable. This misalignment could stem from problems in the photometry of that pixel, such as contamination from other objects or improper background subtraction. However, it is more likely due to assumptions in the SPS models, including the prior ranges. On the other hand, we observe for some of the pixels there multiple possible SED templates and the model correctly interprets this, so that the posterior distributions are wider, especially at low S/N and at high redshifts where several filters are only fitted with upper limits. The best example for this is the pixel outside $1R_{\rm{eff}}$ of 211273. Finally, a few posterior samples can produce SEDs that appear inconsistent with the photometry, as seen for example in galaxy 188081. This object is very compact and dominated by emission lines, making it particularly difficult to model within our simulation framework, which uses Cloudy from FSPS without emission line fine-tuning. The apparent deviations mainly stem from imperfect modelling of emission lines in the redder part of the spectrum. These cases highlight the non-linear relationship between physical parameters and flux space, where modest parameter shifts can lead to significant SED variation.
We include in the same figure the best fit of Prospector in pink, as well as the Synthesizer fit in green. While the fit of Prospector uses the same SPS modelling assumptions, and the deviations we see with respect of our model are mainly random, the differences with Synthesizer are systematic, as we can clearly see different spectral templates. This is more evident for high-redshift galaxies and show the importance of the choice of libraries and parameters for stellar population properties estimation and of marginalising over all the different templates to properly capture the true uncertainties.
@@ Appendix F Profiles of stellar population properties
We compute profiles of stellar mass density, mass-weighted age and dust attenuation, from posterior distributions obtained with the parametric model. In Fig. 15, we show in blue the profiles of the properties computed from elliptical apertures with the semimajor axis, the ellipticity and the position angle obtained from the JADES photometry catalogue (Eisenstein et al. 2023a), and the centres fixed as the pixel with higher stellar masses (medians of the posteriors). The effective radius we use to normalise the distances to the centre was also obtained from the JADES catalogue. The profiles correspond to the medians of the posterior distributions and the errors to the $1\sigma$ level, normalised to the values in the centre of the galaxy. It is important to note again that the profiles below the PSF FWHM of F [^4] W are mainly flat, and after this radial scale vary significantly.
We recover what we see in the previous maps, the stellar mass density profiles are smooth and do not suffer strong variations, while the other properties, especifically for the first three galaxies where we reach larger radius, show important variations. While the uncertainties ($1\sigma$) are considerable, we find similar profiles for both three low redshift galaxies (first row), with medium-age centres, old regions between 0.20.50.2-0.5 $R_{\rm{eff}}$ and younger outskirts. For the high-redshift galaxies, we obtain flatter profiles, specially for 118081 where we are mostly inside the FHWM of F [^4] W due to the compactness of the galaxy. For the first four galaxies the centres are more dust attenuated than the outskirts, but both 118081 and 206146 have very little dust contribution in the centre, compatible with the high redshift of these sources.
@@ Appendix G Offset in stellar mass estimations with different priors for the SFHs
For the $\tau$-delayed model applied to the Dirichlet simulation we observe a significant average underestimation of $0.15$ dex. This is expected because the fixed form of the $\tau$-delayed prior is not able to capture early episodes of star formation (e.g. Carnall et al. 2019; Lower et al. 2020). We also compute the fraction of the true total stellar mass formed before the median $t_{i}$ obtained with the $\tau$-delayed model, according to the simulated Dirichlet SFH. We find a correlation between the residuals and this fraction, showing that the largest underestimations of total stellar mass for the $\tau$-delayed model actually occur when there are episodes of star formation that contribute significantly to the mass, but are obscured by more moderate and recent bursts of star formation.
@@
In the second row, the simulation is performed with a $\tau$-delayed prior. While in the left case, where $\tau$ is large and thus the SFH is extended, the Dirichlet model performs well and also obtains a recent SFR that is close to the true value, in the right case the Dirichlet model predicts previous episodes of star formation that are not realistic, leading to $69\%$ of the mass predicted to form before the star formation actually starts. Nevertheless, in this case we observe that some of the SFHs drawn from the posterior samples are fully compatible with the simulation. This flexibility of the Dirichlet prior, perhaps excessive, allows us to marginalise through a complete set of behaviours, and in this case the large uncertainties in constraining the SFHs are present in the posterior distributions of all the parameters. This leads to broader posterior distributions, as already shown in Fig. 12, which are more representative of the ignorance of the past SFHs of true galaxies. This effect is also found in the pixel-by-pixel analysis of true JWST galaxies. While for most galaxies the flexibility of the pixel-by-pixel analysis probably does not require complex SFHs at the pixel level, one pixel may represent several SSPs with non-trivial SFHs, especially at high redshift. Ultimately, we should look for physically motivated priors that match the burstiness and star formation scales of the real Universe. This approach, already explored in the works of Caplar & Tacchella (2019); Tacchella et al. (2020); Iyer et al. (2024); Wan et al. (2024), and Wang et al. (2025), for instance, is crucial to properly assess the estimation of stellar population parameters with well-calibrated uncertainties from photometry alone.